QUESTION 16 Miller Manufacturings degree of operating leverage is 1.5. Warren Corporations degree of operating leverage is
Answer #1

Answer is 2 times

The degree of operating leverage is the leverage what a firm gets out of its fixed cost component in the total cost. The higher the fixed cost and lower variable cost the degree of operating leverage is higher. The higher the operating leverage the profit growth or decline will be higher since increase or decrease in net income is calculated by sales growth multiplied by degree of operating leverage.

Warren’s degree of operating leverage 4.5

Miller Manufacturing degree of operating leverage is 1.5

So Warren’s earnings will increase or decrease 2 times over Miller’s equal increase or decrease in sales.

Calculation (4.5/1.5)= 3-1 = 2 times